# Non-degenerate mixed maps and contact structures

### v1

@@ 4. Natural and Milnor fillable structures
(4.1) $\frac{{\bf i}}{2}d(\Theta\circ\psi)_{z}(R)$ $=\sum_{i,j}\alpha_{j}\left(\frac{\partial\operatorname{Re}(\psi_{i})}{\partial z_{j}}-\frac{\partial\operatorname{Re}(\psi_{i})}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right)-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{j}\left(\frac{\partial\operatorname{Im}(\psi_{i})}{\partial z_{j}}-\frac{\partial\operatorname{Im}(\psi_{i})}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right)$
$=\sum_{i,j}\alpha_{j}\frac{\partial\operatorname{Re}(\psi_{i})}{\partial y_{j}}-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{j}\frac{\partial\operatorname{Im}(\psi_{i})}{\partial y_{j}}.$
Let $h=z_{1}$ and set $r_{1}=\sum_{i}\frac{\partial\operatorname{Re}\psi_{i}}{\partial y_{1}}$ and $r_{2}=\sum_{i}\frac{\partial\operatorname{Im}\psi_{i}}{\partial y_{1}}$. By the local form of immersions, we may suppose the Jacobian matrix of $\psi$ in the coordinates $(x,y)$ has the following local form:
$\begin{pmatrix}(D\psi)_{2n-(2k+1)\times 2n-(2k+1)}&{0}_{2n-(2k+1)\times 2k+1}\\ {0}_{(2k+1)\times 2n-(2k+1)}&{0}_{2k+1\times 2k+1}\end{pmatrix}$,
where $D\psi$ is an invertible square matrix. If (4.1) is zero at a point $p$ and $\Theta$ is defined by $h$, then $r_{1}-r_{2}=0$. Multiplying the rows corresponding to the imaginary parts and summing up, we obtain a null column and this leads to a contradiction with the fact that $\psi$ is a diffeomorphism on $K_{V}$. ∎
For the next results, we suppose the link of the mixed map $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is ambient Milnor fillable.
@@
Suppose that $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is strictly holomorphic-like. If the restriction of $\xi_{r}$ to the binding $\phi(N)$ is a contact structure, then the induced Milnor fillable and the natural contact structures are isotopic.
@@
In this case, $\xi_{r}$ and $\xi_{G}$ are both adapted to $(\psi^{*}(\Theta),\phi(N))$. Consider a fiber of $\psi^{*}\Theta$ and note that it is endowed with two symplectomorphic structures, namely, $\psi^{*}(d\alpha)$ and $d\alpha$. Besides being isotopic, the symplectic structures on the completions are also symplectomorphic, and thus the result follows from [11, Proposition 9]. ∎
@@
Let $G:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed Hamm ICIS germ given as the pullback $G=\phi_{a,b}^{*}H$ of some complex map germ $H:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ by a homogeneous mixed covering $\phi_{a,b}$, where $a>b$. Then the induced Milnor fillable structure is isotopic with the natural contact structure.
@@
The Milnor fillable contact structure is induced from the diffeomorphism $\psi:=\psi_{1}$ of Theorem 2.12 and the binding $N=K_{F}\cap\{z_{1}=0\}$ is mapped to $\psi(N)=K_{G}\cap\{z_{1}=0\}$, where $F$ is the associated Hamm ICIS. But $\psi(N)$ is the link of $G^{I}$, where $I=\{2,\dots,n\}$. Furthermore, $G^{I}$ is a strictly holomorphic-like mixed function and so $\psi(N)$ is a contact submanifold of $\mathbb{S}_{r}^{2n-3}$. The result follows from Corollary 4.2. ∎
@@ Acknowledgments
We thank Patrick Popescu-Pampu for useful discussions and for calling to our attention the fact remarked in Proposition 3.11. We also thank Maria Aparecida Ruas, Cidinha, for helpful conversations.

### v3

@@ 4. Natural and Milnor fillable structures
$d(\Theta\circ\psi)_{z}(R)=\sum_{i}$ $\frac{1}{z_{1}}\left[\frac{\partial}{\partial z_{i}}\left(\frac{\psi_{1}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{1}}{2}\right)z_{i}-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\left(\frac{\psi_{1}+{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{1}}{2}\right){}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\right]$
- $\frac{1}{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}\left[\frac{\partial}{\partial z_{i}}\left(\frac{\psi_{1}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{1}}{2{\bf i}}\right)z_{i}-\frac{\partial}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\left(\frac{\psi_{1}-{}\mkern 3.0mu\overline{\mkern-3.0mu\psi}_{1}}{2{\bf i}}\right){}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\right].$
By the local form of immersions, we may suppose that $\psi$ is locally given by:
$\psi(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=(z_{1},z_{2},\dots,\operatorname{Re}z_{l},0,\dots,0),$
where l=nkl=n-k, since the (real) dimension of $K_{G}$ is 2n2k12n-2k-1. Suppose that there exists $p\in K_{G}\setminus h^{-1}(0)$ such that $d(\Theta\circ\psi)_{z}(R)=0$ at $p$. The equation above reduces to:
$\frac{1}{z_{1}}\left[\frac{z_{1}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}{2}\right]=\frac{1}{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}\left[\frac{z_{1}+{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}}{2{\bf i}}\right].$
This gives $(\operatorname{Im}z_{1}){}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}=\operatorname{Re}(z_{1})z_{1}$, and thus $z_{1}=0$, which leads to a contradiction. ∎
It is worth remarking that a complex holomorphic function in one single coordinate, such as $h$ in the above proof, was also used in [41] to construct open books adapted to the natural contact structure on complex Pham-Brieskorn manifolds, a result generalized by [6, Theorem 3.9]. For the next results, we suppose the link of the mixed map $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is ambient Milnor fillable.
@@
Suppose that $G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is strictly holomorphic-like. If the restriction of $\xi_{r}$ to the binding $\phi(N)$ is a positive contact structure, then the induced Milnor fillable and the natural contact structures are isotopic.
@@
In this case, $\xi_{r}$ and $\xi_{G}$ are both adapted to $(\psi^{*}(\Theta),\phi(N))$. Consider a fiber of $\psi^{*}\Theta$ and note that it is endowed with two symplectomorphic structures, namely, $\psi^{*}(d\alpha)$ and $d\alpha$. Besides being isotopic, the symplectic structures on the completions are also symplectomorphic, and thus the result follows from [13, Proposition 9]. ∎
@@
Consider the family $G_{t}$ defined in (2.4) of mixed map germs. Observe that, for each $t\in 0<t\leq 1$ fixed, the Newton polyhedron of $G_{t}$ is the same as that of the complex holomorphic Hamm map $F$.
@@
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a complex Hamm map and $G_{t}$ a deformation as in (2.4). Then the induced Milnor fillable structure is isotopic with the natural contact structure on the link $K_{G_{t}}$ of $G_{t}$ for each $0<t\leq 1$.
@@
The Milnor fillable contact structure is induced from the inverse diffeomorphism $\psi_{t}^{-1}$ of Theorem 2.24 and the binding $N=K_{F}\cap\{z_{1}=0\}$ is mapped to $\psi_{t}(N)=K_{G_{t}}\cap\{z_{1}=0\}$. But $\psi(N)$ is the link of $G^{I}_{t}$, where $I=\{2,\dots,n\}$. Furthermore, $G^{I}_{t}$ is a strictly holomorphic-like mixed function by Proposition 3.13 and Theorem 3.14, and thus $\psi_{t}(N)$ is a contact submanifold of $\mathbb{S}_{r}^{2n-3}$. The result follows from Corollary 4.2. ∎
@@ Acknowledgments
We thank Patrick Popescu-Pampu for useful discussions and for calling to our attention the fact remarked in Proposition 3.11. We are also grateful to Maria Aparecida Ruas (in memoriam), Cidinha, for valuable conversations and for her inspiring example as both a human being and a mathematician.